# Discovering and exploiting active sensing motifs for estimation

### v1

@@ Abstract
From organisms to machines, autonomous systems rely on measured sensory cues to estimate unknown information about themselves or their environment. For nonlinear systems, carefully selected sensor motion can be exploited to extract information that is otherwise unavailable, i.e. active sensing. Empirical, yet mathematically rigorous, tools are needed to (1) quantify how sensor movement can contribute to estimation performance, and (2) leverage this knowledge to improve state estimates. Here, we introduce “BOUNDS: Bounding Observability for Uncertain Nonlinear Dynamic Systems”, and Python package pybounds, which can discover patterns of sensor motion that increase information for individual state variables. Crucially, it is suitable for partially observable nonlinear systems, accounts for sensor noise, and can be applied to either simulated or observed trajectories. We demonstrate BOUNDS through a case study on a flying agent with limited sensors, showing how active sensing can be leveraged to estimate key variables such as ground speed, altitude, and ambient wind direction. Finally, we present a framework to refine sporadic estimates from bouts of active sensing that combines data-driven state and observability estimation from artificial neural networks with model-based estimation, which we call the Augmented Information Kalman Filter (AI-KF). We validate our framework using altitude estimation given GPS-denied data from an outdoor quadcopter flight. Collectively, our work will help decode active sensing strategies and inform the design of estimation algorithms in sensorimotor systems.
@@
observability; Fisher information; active sensing; anemosensing; quadcopter
@@ Introduction
A fundamental challenge for any autonomous agent—organism or robotic system—is to combine and transform incoming sensory cues into estimates of task-relevant variables. For example, pixels measured by a camera over time can be transformed into information about self-motion. In many cases, variables of interest are not linearly related to available sensor measurements or their derivatives, which complicates estimation. Optic flow cues, for instance, encode a nonlinear function of velocity and distance, making decoding either of these two quantities independently a difficult task Martinelli2012, VanBreugel2014.
For a moving agent with nonlinear dynamics, sensory cues are shaped by self-motion or independently moving sensors. Therefore, strategically applied sensor motion can help decouple correlated variables and enhance an agent’s ability to perform an estimation task, a process we refer to as active sensing. For example, autonomous agents that want to estimate their velocity or distance relative to nearby objects from optic flow can not do so from a stationary camera. Optic flow captures the ratio of velocity-to-distance ($r=v/z$), making it difficult to estimate either one independently. However, by changing velocity with known acceleration, optic flow can be decoupled to estimate both distance VanBreugel2014a and velocity Lingenfelter2021 via the chain rule (i.e. $\dot{r}=\dot{v}/z-r\dot{z}$ can be solved for $z$). In this example, an agent would simply have to accelerate its vision system to apply this basic active sensing principle—a behavior exhibited by animals such as mice Parker2022 and mantids Poteser1995.
A major barrier to applying active sensing principles is understanding which patterns of sensor movements are most appropriate for decoupling key variables. How can we understand what extra information can be acquired via sensor movements? Existing approaches leverage the dynamical systems concept observability, which describes the ability to infer the state variables of a system from the accessible measurements and inputs hermann2003nonlinear. Analytical tools exist for making a binary assessment of nonlinear observability of individual state variables VanBreugel2021a, Anguelova2004. Quantifying observability levels typically relies on measures based on the singular- or eigen- values of the observability matrix or its Gramian marques1986relative, rui2008observable, Krener2009, which captures the observability of state combinations. Despite recent efforts Cellini2023, existing methods for quantifying observability levels remain poorly suited for isolating individual state observability for high dimensional and/or partially observable nonlinear systems (i.e. systems for which some state variables cannot be estimated), do not offer a physical unit for observability measures, and do not consider the impact of sensor noise. Furthermore, while data-driven observability tools exist Massiani2024, there are no established approaches for applying observability tools to experimental data, such as a measured agent trajectory.
Here we introduce a novel computational pipeline—“BOUNDS: Bounding Observability for Uncertain Nonlinear Dynamic Systems” [^1] Note: an early version of our pipeline appears in a different preprintcellini2024discovering, which was never submitted for publication. The prior preprint will be updated to focus on the application of BOUNDS to understanding wind estimation in flying insects.—that leverages tools from control and information theory to address these shortcomings. Principally, BOUNDS empirically determines the observability levels of individual state variables and computes how observability changes as sensors move. We extend established methods for empirically calculating an observability matrix ($\mathcal{O}$) for nonlinear systems Cellini2023, Lall1999 to work with real trajectory data through a model predictive control framework schwenzer2021review. To take sensor noise properties into account, we calculate the Fisher information matrix ($\mathcal{F}$) associated with $\mathcal{O}$ Berger1985, i.e. the stochastic observability Gramian Boyacoglu2024. We introduce a regularization to invert $\mathcal{F}$ under rank-deficient conditions so that we can apply the Cramér–Rao bound Cramer1946 on partially observable systems and quantify observability with meaningful units. Iterating our pipeline along a dynamic state trajectory intuitively reveals active sensing motifs. Our approach can answer key questions pertaining to sensorimotor systems, such as:
  1. 1.
    Which state variables can be estimated?
  2. 2.
    Which patterns of sensor motion are helpful?
  3. 3.
    Which sensors are required?
BOUNDS can be used as an analysis tool for measured state trajectories or alternatively as a design tool for robotic systems performing state estimation.
Another significant challenge is how to integrate active sensing and observability information for state estimation. It is neither efficient nor practical to constantly perform sensing movements that are unrelated to non-sensing tasks; it would be more efficient to only perform active sensing as needed Biswas2023, Karagoz2025. Classical state estimation methods, however, do not consider time-varying observability corresponding to active sensing, especially not for individual state variables Ramos2021. How might agents continuously compute a state estimate with only sporadic bouts of active sensing that are not corrupted by periods of low observability?
For nonlinear and/or partially observable systems, model based estimators—for instance, the ubiquitous Kalman filter (KF) framework urrea2021kalman—can exhibit instability and an increased sensitivity to noise or initialization in weakly observable dimensions Butcher2017, Bocquet2017, Devos2021. Model-based filters typically rely on an arbitrary initialization “guess” for the state and state covariance that is then incremented based on incoming measurement/inputs. While this approach works well—and will eventually converge to an optimal filter for observable, linear dynamics—there is no guarantee that these filters (including nonlinear extensions EKF and UKF julier1997new, wan2000unscented, Wan2006) will converge for systems that significantly deviate from linearity and observability assumptions. The critical flaw of iterative estimation schemes applied to nonlinear systems, such as the KF, is that measurement information is only correctly incorporated when the state estimate itself is already accurate. When prior state estimates are inaccurate, information is not efficiently extracted from the measurements, and is effectively lost. Data-driven estimators may be able to circumvent this drawback by accessing an arbitrarily large time window of measurements to directly estimate observable state variables. Such estimators, however, will produce very inaccurate estimates during periods of low observability.
We present a novel state estimation framework that leverages the observability information generated by BOUNDS to stitch together sporadic estimates corresponding to bouts of active sensing. Our framework blends two widely used estimation strategies: the Kalman filter and data-driven feed-forward artificial neural network (ANN) estimators. To combine these two estimators in a statistically meaningful manner, we leverage our BOUNDS framework to learn and assign an estimated variance to the ANN estimates and incorporate this estimate as an augmented measurement in the Kalman filter. We call this framework the Augmented Information Kalman Filter (AI-KF) because it augments the traditional KF with information derived from a recent time history of measurements. Although our approach has broad applications, we illustrate its utility through a specific active sensing challenge: how a flying agent with limited sensors could employ active movements to estimate key variables pertaining to it’s own movement (altitude, ground speed) and external variables (ambient wind direction). We show that our AI-KF framework outperforms a traditional Unscented Kalman Filter (UKF) in scenarios with poor initial state or covariance guesses, unmodeled disturbances, and trajectories with brief periods of observability.
@@
Figure 1: Overview of BOUNDS, a method for discovering active sensing motifs with empirical observability. a. A simulated (or measured) state trajectory (in this case, of a flying agent) that is described by a collection of time-series of observed state variables $\mathbf{\tilde{X}}$. b. Model predictive control (MPC) is used to find the time-series of control inputs $\mathbf{U}$ required to reconstruct the state trajectory in a through a simulation model. c. The MPC control inputs $\mathbf{U}$ are used to simulate the dynamics associated with the state trajectory, yielding the nominal reconstructed trajectory $\mathbf{X}$. A measurement model $\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$ simulates the sensor measurements. d. Each initial state variable $x_{i,0}$ is perturbed in the positive $x_{i,0}^{i+}$ and negative $x_{i,0}^{i-}$ directions by $\epsilon$. Then the corresponding measurements over time, $\mathbf{Y}^{i+}$ and $\mathbf{Y}^{i-}$, are computed. The difference between these measurements is used to calculate a numerical Jacobian: $\Delta\mathbf{Y}^{i}/2\epsilon$, yielding one column of the observability matrix $\mathcal{O}$. e. An illustration of an observability matrix $\mathcal{O}$, where the color grid indicates the values in the matrix (corresponding to d). The columns indicate the difference of the initial state perturbation $\Delta x_{i,0}$ and the rows indicate the corresponding difference in measurements $\Delta\mathbf{y}_{k}$ over time for a discrete number of time steps $\omega$. f. An illustration of the Fisher information matrix, calculated as $\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O}$, where $\mathcal{R}$ is the block diagonal sensor noise covariance matrix. Colors indicate matrix values as in e. g. The inverse of the Fisher information matrix $\mathcal{F}^{-1}$, which encodes the minimum error covariance of the estimate of the state’s initial condition $\mathbf{x}_{0}$. The diagonal elements represent the minimum error variance of each individual state variable—inversely correlated with the observability level—where smaller values indicate a higher observability level. h. The minimum error variance time-series (normalized units) of all state variables (diagonal of $\mathcal{F}^{-1}$) over time for the first three sensors $y_{1},y_{2},y_{3}]$ (left) and sensors $y_{2},\cdots,y_{p}]$ (right). Note that some states are persistently observable or unobservable, while other states exhibit time-varying observability, which is dependent on the sensor set. i. The minimum error variance time-series of the first state variable $x_{1}$ for the first three sensors $y_{1},y_{2},y_{3}]$ (left) and the corresponding correlation with the active sensing motifs in the simulated trajectory (right) from a. Time-series corresponds to the first column in h. (left). By correlating the time-series with the trajectory, we reveal that only turns in heading improve the observability of $x_{1}$. To make the relationship between temporal patterns in observability and the features of the trajectory more intuitive, we shift the time series back in time by half of the sliding window length, $\omega/2$. j. Same as i but for state variable $x_{2}$ and sensors $y_{2},\cdots,y_{p}]$. Corresponds to the second column in h (right). This sensor set renders acceleration/deceleration and turns offset to heading observable, but headings turns have no effect.
@@ Results
To describe our computational pipeline, as well as new scientific discoveries, we have organized this section as follows. First, we provide a step-by-step description of the BOUNDS computational pipeline using Figure 1 as a visual organization of these steps. Second, we introduce a specific case study—state estimation for a flying agent—to showcase the different ways our method can be used to discover active sensing motifs from simulations and data. Finally, we introduce a framework for incorporating knowledge of observability into an estimator to provide continuous estimates across sporadic bouts of active sensing, and again illustrate this approach using our case study.
@@ 1. Determine state variables and collect state trajectory time-series
2. Define a model
$\mathbf{x}_{k+1}$ $=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k}),$ (1)
where $\mathbf{x}_{k}$ is the true state vector representing $n$ state variables, $\mathbf{u}_{k}$ is the input vector, and $\mathbf{\tilde{y}}_{k}$ is the measurement vector corresponding to $p$ measurements at time step $k$. The function $\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})$ describes the system dynamics. The function $\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$ describes the measurement dynamics, and $\mathbf{v}_{k}$ is zero-mean additive Gaussian noise with covariance matrix $R_{k}$. The effect of the noise on observability is taken care of in Step 5. Steps 3-4 only use the deterministic measurement equation $\mathbf{y}_{k}=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$. A benefit of our empirical approach is that the model does not have to be closed-form, and can instead be a simulation, computational, or data-driven model Singh2005, Krener2009—for instance, a physics engine model of Drosophila Vaxenburg2024 or finite element model of a flapping wing Boyacoglu2024a. The only requirement is the ability to simulate the system forward in time.
3. Reconstruct state trajectory through model predictive control
Next, our method applies model predictive control (MPC) Camacho2007 to find the inputs $\mathbf{U}$ required to reconstruct the state trajectory with the chosen model (LABEL:fig:figure_1_b). This step is critical for nonlinear or time-varying models where observability can change as a function of the state and input values. The MPC determined inputs are used to simulate the nominal trajectory $\mathbf{X}$ (which should be very similar to the observed trajectory $\mathbf{\tilde{X}}$) and associated measurements $\mathbf{Y}$ (LABEL:fig:figure_1_c).
4. Construct observability matrix
To quantify the observability of individual state variables, we next compute the observability matrix $\mathcal{O}$ empirically. $\mathcal{O}$ principally encodes information about the sensitivity of the sensor measurements $\mathbf{y}_{k}$ to small changes in an initial state $\mathbf{x}_{0}$. It can be compactly represented in discrete-time as the Jacobian of the measurements over time with respect to the initial state, $\Delta\mathbf{Y}_{\omega}/\Delta\mathbf{x}_{0}$ Georges2020. Here we use $\mathbf{Y}_{\omega}$ to denote the collection of $p$ measurements for each time step from k=[0:w1]k=[0:w-1]. Following established methodsSingh2005, Krener2009, we empirically approximate the Jacobian by applying small perturbations to each initial state variable $x_{i,0}$, one at a time, in the positive ${x}^{+}_{i,0}$ and negative ${x}^{-}_{i,0}$ directions. We simulate the system using the inputs $\mathbf{u}_{k}$ calculated in step 3 (they are not recalculated for the perturbed trajectory), and then compute the difference in the corresponding change in sensory measurements over time $\Delta\mathbf{Y}_{\omega}^{i}=\mathbf{Y}_{\omega}^{i+}-\mathbf{Y}_{\omega}^{i-}$ (LABEL:fig:figure_1_d). The end result for each state perturbation is a column vector of length $p\times\omega$, corresponding to $p$ measurements over $\omega$ time steps. Repeating these steps for each state variable and stacking the resulting column vectors side-by-side yields the full empirically determined Jacobian, $\Delta\mathbf{Y}_{\omega}/\Delta\mathbf{x}_{0}$, which is equivalent to the empirical observability matrix $\mathcal{O}$ (LABEL:fig:figure_1_e).
5. Compute the Fisher information matrix and its inverse
To compute observability levels for individual state variables in physically interpretable units, we construct the Fisher information matrix as
$\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O},$ (2)
where $\mathcal{R}$ is a block diagonal matrix of the measurement noise covariance matrices at each time step. For the deterministic dynamics we consider, the process noise $Q$ is 0 (see Methods) (LABEL:fig:figure_1_f) Hwang1985, Jauffret2007. The Fisher information matrix captures the amount of information that a known random variable, in our case the measurements $\mathbf{Y}_{w}$, carries about unknown parameters, in our case the initial state $\mathbf{x}_{0}$. The Cramér–Rao bound states that the inverse of the Fisher information matrix $\mathcal{F}^{-1}$ sets the lower bound on the error covariance of the state estimate Cramer1946. Thus, the diagonal elements of $\mathcal{F}^{-1}$ tell us the minimum possible variance in the error for each state variable (for any estimator), inversely correlating with observability (LABEL:fig:figure_1_g).
In cases where there is at least one unobservable state variable, $\mathcal{F}$ will not be invertible. Therefore, to compute $\mathcal{F}^{-1}$ we add a small regularization term to ensure invertibility chernoff1953locally:
For unobservable state variables, this “Chernoff inverse” yields infinity in the corresponding diagonal elements, whereas the more traditional Moore-Penrose inverse Penrose1955 yields misleading values. For computational efficiency, we use a small real value for $\lambda$, which places an upper bound on the minimum error covariance. Thus, $\lambda$ must be chosen carefully, so as not to inject too much “artificial” observability into the system, while still allowing an inverse to be computed (see Methods).
6. Perform coordinate transformation
The coordinates of both $\mathcal{O}$ and $\mathcal{F}$ are principally defined in the state space $\mathcal{X}$, which means that $\mathcal{F}^{-1}$ only provides information with respect to the individual state variables in $\mathbf{x}$. However, in many cases, it is beneficial to evaluate observability in a set of new coordinates $\mathbf{z}$. For instance, a system may be defined in Cartesian coordinates, but the polar representation of the state may be the most relevant for observability analysis. The system itself could be redefined in the new coordinates $\mathbf{f}(\mathbf{z}_{k},\mathbf{u}_{k})$ and the observability analysis pipeline could be repeated. However, this is often tedious for complex systems, not computationally efficient, and is not practical for numerical solvers that rely on a specific coordinate frame. We incorporate a simple method to transform the coordinates of $\mathcal{O}$ (or $\mathcal{F}$) from an arbitrary state space coordinate basis $\mathbf{x}$ to a new basis $\mathbf{z}$ that bypasses these drawbacks and can be appended post hoc after computing $\mathcal{O}$ with respect to $\mathbf{x}$.
First, the new basis $\mathbf{z}$ is defined as a transformation of $\mathbf{x}$ such that $\mathbf{z}=T(\mathbf{x})$, where $T(\mathbf{x})$ is a diffeomorphism that preserves the original information content in $\mathbf{x}$ VanTrees2001. Because $\mathcal{O}$ fundamentally represents the Jacobian $\frac{\delta\mathbf{Y}}{\delta\mathbf{x}}$ in continuous time or $\frac{\Delta\mathbf{Y}}{\Delta\mathbf{x}}$ in discrete time, we can compute a coordinate transformation as the Jacobian of $\mathbf{x}$ with respect to $\mathbf{z}$, evaluated at the initial state $\mathbf{x}_{0}$
$\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{x}=\mathbf{x}_{0}}=\left.\left(\frac{\partial T(\mathbf{x})}{\partial\mathbf{x}}\right)^{-1}\right|_{\mathbf{x}=\mathbf{x}_{0}}$ (4)
and transform the coordinates as
$\mathcal{O}_{\mathbf{z}}=\left.\frac{\partial\mathbf{Y}}{\partial\mathbf{z}}\right|_{\mathbf{x}=\mathbf{x}_{0}}=\frac{\partial\mathbf{Y}}{\partial\mathbf{x}}\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{x}=\mathbf{x}_{0}}=\mathcal{O}_{\mathbf{x}}\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{x}=\mathbf{x}_{0}}.$ (5)
In practice, $J_{\mathbf{T}}$ can be computed numerically or symbolically and evaluated at the initial state $\mathbf{x}_{0}$ considered by $\mathcal{O}$. In some cases, it may be necessary to perform multiple coordinate transforms to evaluate the observability of new state variables that may not jointly satisfy the diffeomorphism/information requirement in a single state space.
@@ 7. Reveal active sensing motifs by correlating state dynamics with observability
To illustrate the applications of BOUNDS for identifying active sensing motifs, we apply it to a 3D model of a quadcopter drone flying in the presence of ambient wind (see Methods). The primary state variables consist of the agent’s attitude (roll $\phi$, pitch $\theta$, yaw $\psi$), translational velocity ($v_{x}$, $v_{y}$, $v_{z}$) in the body-level frame, altitude ($z$), and ambient wind speed ($w$) and direction ($\zeta$). We assume the agent has an air flow sensor measures the apparent air flow speed ($a$) and direction ($\gamma$) and an accelerometer that measures the acceleration magnitude ($q$) and direction ($\nu$). We also assume the agent has a vision system that can measure ventral optic flow magnitude ($r=g/z$) and direction ($\beta$), where the ground speed magnitude is defined as $g=\sqrt{{v_{x}}^{2}+{v_{y}}^{2}}$ . In addition, we assume that our agent has access to its own yaw measurement $\psi$. These measurements make up different components of the wind vector “triangle” (LABEL:fig:figure_2_aLABEL:fig:figure_2_b) experienced by any flying agent VanBreugel2021a, and pose interesting active sensing questions that can be answered via BOUNDS.
In general, active sensing motifs will correspond to different combinations of state values, and changes in states, that can be reached via control within the state space $\mathbf{x}$. Therefore, it is best practice to employ a methodical exploration of the state space and identify the regions corresponding to increasing observability. However, for illustrative purposes, we examine a limited set of motifs that can easily be visually identified for a flying agent: 1) acceleration, 2) deceleration, 3) turning while changing heading, 4) turning without changing heading (offset turn), and 5) a small turn through the upwind direction ($\zeta=0$) (LABEL:fig:figure_2_c). We then asked which states become more or less observable for these motifs depending on what sensors were available.
We simulated a state trajectory with a sequential combination of these motifs for constant wind and evaluated the observability of each state variable as a function of the trajectory (LABEL:fig:figure_2_c). For brevity, we focus on the wind direction $\zeta$, altitude $z$, and ground speed $g$ state variables. Note that $g$ is not in the original state space of our model (see Methods) and required our coordinate transformation step in BOUNDS.
@@
We first considered a sensor set consisting of only angular measurements (heading $\psi$, course direction $\beta$, apparent airflow angle $\gamma$). We discovered that only the ambient wind direction state variable could be estimated with these measurements, however a change in heading motif was required and straight flight was not sufficient (LABEL:fig:figure_2_ci). This result is consistent with both analytical observability results VanBreugel2021a and conclusions from optimal estimator designs VanBreugel2022. A novel finding from our analysis was that larger turns generally led to a higher degree of observability but a small turn passing through the upwind direction was also adequate for high observability (LABEL:fig:figure_2_ci). Acceleration/deceleration and offset turn motifs did not markedly increase observability for this sensor set.

### v2

@@ Abstract
From organisms to machines, autonomous systems rely on measured sensory cues to estimate unknown information about themselves or their environment. For nonlinear systems, strategic sensor motion can be leveraged to extract otherwise inaccessible information. This principle, known as active sensing, is widespread in biology yet difficult to study, and remains underutilized in engineered systems due to the challenge of systematically designing active sensing motifs. Here, we introduce the method “BOUNDS: Bounding Observability for Uncertain Nonlinear Dynamic Systems”, and Python package pybounds, which can discover movement motifs that increase the information encoded in sensory cues. To exploit sporadic estimates from bouts of active sensing, we further introduce the Augmented Information Kalman Filter (AI-KF). The AI-KF uses insight from BOUNDS to dynamically fuse neural network and model-based estimation. We demonstrate BOUNDS and the AI-KF on a flying agent model and experimental GPS-denied data from a quadcopter, revealing how specific active movements improve estimates of ground speed, altitude, and wind direction. Altogether, our work will prove useful for designing sensor-minimal autonomous systems and investigating active sensing in living organisms.
@@
active sensing; observability; Fisher information; estimation
@@ Introduction
A fundamental challenge for autonomous agents—organisms 23 or machines 54—is to combine sensory cues into estimates of task-relevant variables 65. This challenge intensifies when variables of interest are nonlinearly related to sensor measurements 50. For example, optic flow encodes the ratio of velocity-to-distance, making it impossible to estimate either quantity independently from any single moment 37, 57. For moving agents with nonlinear dynamics or measurements, strategically applied movement can decouple correlated variables and enhance estimation, a process we refer to as active sensing 66. By accelerating their vision systems, animals like mice 41 and mantids 44 can gauge distance, while visual-inertial sensor fusion in robots requires similar motion to estimate position 6, 45.
For complex estimation tasks, a major barrier to designing active sensing strategies for machines, and understanding their role in organisms, is determining which movements decouple key variables. Existing approaches leverage the control-theoretic concept observability. Closely related to Fisher information 8, observability describes how well system states can be inferred from measurements and inputs 26. Analytic observability tools have yielded qualitative insight into how movement motifs in fish 52, 33, flies 60, and machines 27 can aid in estimation, but their inability to distinguish weak vs. strong observability levels limits their utility, especially in data-driven applications. Current quantitative methods, on the other hand, rely on eigenvalues of the observability Gramian 36, 31, making it difficult to evaluate observability of individual state variables for partially observable nonlinear systems 14. Neither accounts for unknown system inputs, preventing their application to experimental trajectory data. Consequently, no existing approach can quantify observability for individual state variables in partially observable nonlinear systems from experimental data, precisely what is needed to systematically discover active sensing motifs for real world applications.
A second challenge is integrating active sensing into state estimation. Classical estimators like the Kalman filter (KF) framework 56 do not explicitly account for time-varying observability, and can exhibit instability in weakly observable dimensions 11, 5. The critical flaw of iterative schemes like the KF is that measurement information is only correctly incorporated when state estimates are already accurate. When estimates are poor, information is effectively lost. Data-driven approaches that can access a history of information all at once offer a potential resolution to this challenge. Increasing interest in artificial neural networks (ANNs) has spawned numerous ANN-KF hybrids 2 that replace KF components 34, 18, 47, entire filters 25, 32, or precondition inputs 35, 64. None, however, provide a theoretically grounded approach for determining when and how much to weigh model-based predictions with data-driven estimates that draw on a time history of measurements.
Here we introduce BOUNDS (Bounding Observability for Uncertain Nonlinear Dynamic Systems) [^1] An early version appears in 13, which was never submitted for publication., an empirical computational pipeline that quantifies observability of individual state variables along dynamic trajectories while accounting for sensor noise through Fisher information and the Cramér-Rao bound 19. BOUNDS reveals which states are estimable, which movement motifs are helpful, and which sensors are required—both for measured trajectories and robotic system design. We further present the Augmented Information Kalman Filter (AI-KF), which provides a framework for dealing with sporadic estimates from active sensing systems. The AI-KF combines model-based filtering with ANN-based estimation by using BOUNDS-derived variance estimates to merge sporadic active sensing information in a statistically meaningful manner. We demonstrate this framework on a flying agent estimating altitude, ground speed, and wind direction with limited sensors, showing the AI-KF outperforms traditional approaches under poor initialization, unmodeled disturbances, and intermittent observability.
@@
Figure 1: Overview of BOUNDS, a method for discovering active sensing motifs with empirical observability. See Methods for a detailed step-by-step description. a. A simulated (or measured) state trajectory (in this case, of a flying agent) that is described by a collection of time-series of observed state variables $\mathbf{\tilde{X}}$. b. Model predictive control (MPC) is used to find the time-series of control inputs $\mathbf{U}$ required to reconstruct the state trajectory in a through a simulation model. c. The MPC control inputs $\mathbf{U}$ are used to simulate the dynamics associated with the state trajectory, yielding the nominal reconstructed trajectory $\mathbf{X}$. A measurement model $\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$ simulates the sensor measurements. d. Each initial state variable $x_{i,0}$ is perturbed in the positive $x_{i,0}^{i+}$ and negative $x_{i,0}^{i-}$ directions by $\epsilon$. Then the corresponding measurements over time, $\mathbf{Y}^{i+}$ and $\mathbf{Y}^{i-}$, are computed. The difference between these measurements is used to calculate a numerical Jacobian: $\Delta\mathbf{Y}^{i}/2\epsilon$, yielding one column of the observability matrix $\mathcal{O}$. e. An illustration of an observability matrix $\mathcal{O}$, where the color grid indicates the values in the matrix (corresponding to d). The columns indicate the difference of the initial state perturbation $\Delta x_{i,0}$ and the rows indicate the corresponding difference in measurements $\Delta\mathbf{y}_{k}$ over time for a discrete number of time steps $\omega$. f. An illustration of the Fisher information matrix, calculated as $\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O}$, where $\mathcal{R}$ is the block diagonal sensor noise covariance matrix. Colors indicate matrix values as in e. g. The inverse of the Fisher information matrix $\mathcal{F}^{-1}$, which encodes the minimum error covariance of the estimate of the state’s initial condition $\mathbf{x}_{0}$. The diagonal elements represent the minimum error variance of each individual state variable—inversely correlated with the observability level—where smaller values indicate a higher observability level. h. The minimum error variance time-series (normalized units) of all state variables (diagonal of $\mathcal{F}^{-1}$) over time for the first three sensors $y_{1},y_{2},y_{3}]$ (left) and sensors $y_{2},\cdots,y_{p}]$ (right). Note that some states are persistently observable or unobservable, while other states exhibit time-varying observability, which is dependent on the sensor set. i. The minimum error variance time-series of the first state variable $x_{1}$ for the first three sensors $y_{1},y_{2},y_{3}]$ (left) and the corresponding correlation with the active sensing motifs in the simulated trajectory (right) from a. Time-series corresponds to the first column in h. (left). By correlating the time-series with the trajectory, we reveal that only turns in heading improve the observability of $x_{1}$. To make the relationship between temporal patterns in observability and the features of the trajectory more intuitive, we shift the time series back in time by half of the sliding window length, $\omega/2$. j. Same as i but for state variable $x_{2}$ and sensors $y_{2},\cdots,y_{p}]$. Corresponds to the second column in h (right). This sensor set renders acceleration/deceleration and turns offset to heading observable, but headings turns have no effect.
@@ Results
To build a foundation for methodically exploring how movement can enhance estimation of individual state variables, we developed BOUNDS. BOUNDS begins with two user inputs: a discrete time-series state trajectory $\mathbf{\tilde{X}}$—derived either from simulations or real data—describing agent or sensor movement and environmental variables (LABEL:fig:figure_1_a), and a model defining the dynamics of the agent’s movement ($\mathbf{x_{k+1}}=\mathbf{f}(\mathbf{x_{k}},\mathbf{u_{k}})$) and measurements ($\mathbf{y}=\mathbf{h}(\mathbf{x_{k}},\mathbf{u_{k}})$) (LABEL:fig:figure_1_b). Our empirical approach accommodates any forward-simulable model—closed-form, physics engine, or data-driven. Control inputs can be provided, or in their absence, we apply model predictive control (MPC) to determine the inputs $\mathbf{U}$ that produce the nominal trajectory $\mathbf{X}$, from which we calculate the measurements $\mathbf{Y}$ with the model (LABEL:fig:figure_1_c).
To quantify how much information the measurements provide about each state variable, we determine the sensitivity of the measurements to small perturbations of each state (LABEL:fig:figure_1_d). This is accomplished empirically by perturbing each state variable $x_{i,0}$ (one at a time) in both directions by a small amount $\epsilon$, without changing the control inputs from the nominal values, and recording how the subsequent measurements $\mathbf{Y}$ evolve 49, 31. Repeating this process for each state variable yields the Jacobian $\Delta\mathbf{Y}/\Delta\mathbf{x}_{0}$, which constitutes the empirically determined observability matrix $\mathcal{O}$ 24 (LABEL:fig:figure_1_e).
We then compute the Fisher information matrix $\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O}$ (assuming no process noise)8, where $\mathcal{R}$ accounts for measurement noise covariance (LABEL:fig:figure_1_f). The Cramér–Rao bound states that $\mathcal{F}^{-1}$ provides the lower bound on estimation error variance, but for partially observable systems $\mathcal{F}$ is singular. A key innovation is our regularized inverse: we compute $\mathcal{F}^{-1}=\lim_{\lambda\to 0^{+}}[\mathcal{F}+\lambda I_{n\times n}]^{-1}$. This “Chernoff inverse” 17 yields infinity for unobservable states while providing bounded error estimates for observable ones, unlike the traditional pseudo-inverse 43 that produces misleading values. In practice, we do not evaluate the limit, instead choosing $\lambda$ to be smaller than the largest non-zero singular value of $\mathcal{F}$. The diagonal elements of $\mathcal{F}^{-1}$ thus quantify minimum error variance for each state variable with meaningful physical units (LABEL:fig:figure_1_g).
Calculating $\mathcal{F}^{-1}$ in short sliding windows of length $\omega$ reveals how observability evolves along trajectories (LABEL:fig:figure_1_hLABEL:fig:figure_1_j). (Note that the time window need not be contiguous, see Methods.) In our toy example, changes in heading angle produce order-of-magnitude decreases in minimum error variance for state $x_{1}$ when using sensors $y_{1},y_{2},y_{3}]$ (LABEL:fig:figure_1_i, LABEL:fig:figure_1_h). Conversely, state $x_{2}$ becomes observable during acceleration and offset turns, but requires sensors $y_{2}$ to $y_{p}$ and does not need $y_{1}$ (LABEL:fig:figure_1_j, LABEL:fig:figure_1_h). This illustrates a fundamental insight: each state variable can have distinct observability properties and active sensing motifs. BOUNDS thus reveals which combinations of sensor motion, available measurements, and state variables enable effective estimation.
@@
We applied BOUNDS to a 3D kinematic quadcopter model flying in ambient wind to discover active sensing motifs (see Methods). The primary state variables consist of the agent’s attitude (roll $\phi$, pitch $\theta$, yaw $\psi$), translational velocity ($v_{x}$, $v_{y}$, $v_{z}$) in the body-level frame, altitude ($z$), and ambient wind speed ($w$) and direction ($\zeta$). We assume that the agent has several body-level sensors: heading $\psi$, an air flow sensor to measure apparent air flow speed ($a$) and direction ($\gamma$), an accelerometer that provides the acceleration magnitude ($q$) and direction ($\eta$), and a vision system that measures ventral optic flow magnitude ($r=g/z$) and direction ($\beta$). These measurements make up components of the wind vector “triangle” (LABEL:fig:figure_2_aLABEL:fig:figure_2_b) experienced by any flying agent 60, and pose interesting active sensing questions that can be answered via BOUNDS.
We examined four movement motifs for our agent flying in constant wind: 1) constant course acceleration and deceleration, 2) turning while changing heading, 3) turning without changing heading (offset turn), and 4) a small turn through the upwind direction (LABEL:fig:figure_2_c). We then asked which states became more or less observable for these motifs depending on which sensors are available. For brevity, we focus on the state variables wind direction $\zeta$, altitude $z$, and ground speed $g$. Note that $g$ is not in the original state space of our model (see Methods) and requires a coordinate transformation, which can be done post-hoc by applying the chain rule to transform $\mathcal{O}$ through the Jacobian of the coordinate change (see Methods). A summary of the conditions for observability is given in Table 1, and discussed in the following paragraphs. Our results are not particularly sensitive to the model type: adding the complexity of a full dynamic quadcopter model yielded similar results (Figure S2).
@@
With only angular measurements (heading $\psi$, course direction $\beta$, apparent airflow angle $\gamma$), wind direction alone is observable—but only during heading changes, not straight flight (LABEL:fig:figure_2_ci). This result is consistent with both analytical observability results 60 and conclusions from optimal estimator designs 58. Larger turns generally increased observability, though small turns through the upwind direction sufficed (LABEL:fig:figure_2_ci). Acceleration/deceleration and offset turn motifs did not markedly increase observability for this sensor set.